OpenStudy (anonymous):

14 is 17% of what number? Need the algebraic expression.

OpenStudy (tkhunny):

14 is 17% of (SomeNumber) ==> (Some Number)*0.17 = 14 It's a matter of translation.

OpenStudy (anonymous):

N(.17)=14 ?? or 14 over .17 = .17n over .17 ??

OpenStudy (tkhunny):

Go with courage. When you do things as they shoud be done, don't be discouraged by an odd appearance. N(0.17) = 14 N = 14/0.17 = 82.353